### v1

# Investigating Disordered Granular Matter via Ordered Geometric Fragmentation

Abstract
The evolution of occupied volume under progressive fragmentation of granular matter is studied using a purely geometric model. Rather than modeling disorder directly, properties of disordered granular assemblies are investigated by analyzing an associated family of highly ordered reference configurations that provide sharp upper bounds on accessible volume. Grains are idealized as fragments derived from a hypothetical elongated parent prism with square cross section, sequentially sliced and reassembled into configurations that maximize the enclosed volume.
Analytic expressions are derived for the maximal attainable volume at each fragmentation stage. The volume evolution is non-monotonic: initial fragmentation produces structures whose volume exceeds that of the original object, while further fragmentation leads to monotonic decrease converging to a limiting value of 5/4 times the initial volume, independent of the total number of fragments.
The model reveals pairs of reassembled configurations built from geometrically indistinguishable building blocks yet enclosing different volumes. These conjugate configurations constitute purely geometric analogues of distinct phases connected by rearrangement-induced transitions. Explicit relations link the idealized construction to experimentally measurable grain parameters. Comparison with experimental data on cylindrical rods shows the predicted upper bound falls within the observed packing density range, demonstrating that the model captures essential geometric features despite its simplicity.
I Introduction
Packing problems arise in a wide range of physical contexts, including granular compaction, powder processing, jamming phenomena, and colloidal self-assembly. A substantial body of work has therefore been devoted to understanding how particle shape, anisotropy, and geometric constraints influence packing density and available volume; see, for example, Refs. [1, 2, 3] and references therein. In most studies, the primary objective is to determine optimal or typical packing fractions for prescribed particle shapes within disordered or statistically defined ensembles.
In contrast, the present work addresses a complementary question: how does the volume occupied by granular matter evolve under progressive fragmentation? Rather than modelling disorder directly, this question is examined through a controlled geometric construction based on a single elongated parental object that is sequentially partitioned into identical building blocks. At each fragmentation stage, the maximal enclosed volume attainable after reassembly is evaluated. The resulting configurations are not intended to represent typical granular packings; instead, they provide sharp geometric bounds that elucidate systematic trends in volume evolution induced by fragmentation.
[…]
The packing of elongated particles has been studied extensively both experimentally and theoretically [6, 7, 8, 9]. Systematic experiments have demonstrated that packing fraction decreases with increasing aspect ratio for randomly packed rods [8, 10], with values ranging from $\phi\approx 0.55$ for short rods (aspect ratio $\alpha\approx 4$) to $\phi\approx 0.22$ for highly elongated particles ($\alpha\approx 32$). Importantly, experimental measurements reveal that nominally identical particles under similar preparation conditions can yield a range of packing fractions rather than a single reproducible value [9], indicating the presence of multiple metastable configurations—a phenomenon that finds a natural geometric interpretation in the present model.
While much attention has been devoted to the packing of particles with fixed shapes, the influence of fragmentation on packing density remains less systematically explored. Fragmentation processes — whether mechanical comminution, crushing, or controlled slicing—alter both the size distribution and the geometric properties of constituent grains [11]. The resulting changes in particle aspect ratio and surface morphology can significantly modify packing efficiency. The present work isolates the purely geometric contribution to these effects by considering idealized fragmentation of a single parental object into progressively smaller, identical building blocks.
Building on these ideas, a minimal, fully analytic model is constructed that isolates the geometric consequences of fragmentation and reassembly. Explicit expressions for the maximal enclosed volume at each fragmentation stage are obtained, and several nontrivial structural features are revealed. In particular, the emergence of paired (conjugate) configurations composed of geometrically indistinguishable building blocks, yet enclosing different volumes, is identified — a phenomenon that has no direct analogue in conventional packing problems.
The objectives of this paper are as follows: (i) to quantify the evolution of volume under sequential fragmentation of granular matter using a simple geometric model; (ii) to investigate structural relations among configurations reassembled from differently fragmented building blocks; and (iii) to derive explicit formulae that connect the parameters of the model to experimentally measurable quantities.
II Model and notation
In this work, the geometric properties of granular matter composed of idealised grains are investigated. Friction between grains is assumed to be negligible, and the weight of individual grains is ignored. All grains are assumed to originate from a single, very long hypothetical parental object. The number of grains produced and their individual sizes are determined entirely by the geometry of this initial object.
Grains are modeled as rectangular parallelepipeds with square cross-section. Accordingly, a family of long rectangular prisms (hereafter referred to as initial objects) is considered, denoted by $T_{\,n}^{\,0}$, with square cross-section $a\times a$ and length
$l=2^{\,n+2}a,\qquad n=0,1,2,\dots.$ (1)
(The length is written in this particular form for convenience in subsequent calculations.)
Each initial object may conveniently be viewed as consisting of
$N_{\rm tot}=2^{\,n+2}$
unit cubes of side length $a$. The corresponding volume is
$V_{\,n}^{\,0}=2^{\,n+2}a^{3}.$
Fragmentation of the initial object is carried out stage by stage. At each stage $i=1,2,\dots,n+1$, the object is fragmented according to a rule defined below, and the resulting pieces (grains) are reassembled into a configuration that maximises the volume. The reassembled object is referred to as a tower and is denoted by $T_{\,n}^{\,i}$ , with volume $V_{\,n}^{\,i}$ and relative volume
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}.$
II.1 Fragmentation rule
The fragmentation rule is defined recursively as follows:
  1. 1.
    Stage 0: The initial prism $T_{\,n}^{\,0}$ is defined as a rectangular prism of dimensions $2^{\,n+2}a\times a\times a$.
  2. 2.
    Stage 1: The prism is divided into four equal parts of length $l/4=2^{\,n}a$. These parts are reassembled to maximise the enclosed volume, yielding the tower $T_{\,n}^{\,1}$.
  3. 3.
    Stage 2: Each piece obtained at Stage 1 is divided lengthwise into two equal parts. The resulting eight pieces are reassembled into the tower $T_{\,n}^{\,2}$.
  4. 4.
    This procedure is continued by dividing each piece lengthwise by a factor of two until stage $n+1$, at which point all pieces are unit cubes and the maximal-volume configuration $T_{\,n}^{\,n+1}$ is obtained.
After fragmentation, the elementary building blocks can, in general, be assembled into towers whose cross sections form parallelograms. In the present model, the cross section is restricted to be square. This choice is motivated by the fact that, among all parallelograms with a given perimeter, the square encloses the maximal cross-sectional area.
The square geometry therefore represents an extremal, volume-maximizing configuration within this class and serves as a natural reference case. No claim is made that this construction exhausts all possible configurations, nor that a global maximal-volume theorem is established.
III Fragmentation – reassembly process and volume evolution
Using the fragmentation rule introduced in the previous section, the slicing of long objects is explicitly presented for the first few values of $n$. For each $n$, the sequence of towers $T_{\,n}^{\,i}$ produced at fragmentation stages $i=0,1,\dots,n+1$ is listed together with their volumes $V_{\,n}^{\,i}$ and corresponding relative volumes $R_{\,n}^{\,i}$. These examples motivate the general formulae, which are subsequently derived using mathematical induction.
In the figures below, thick lines indicate grain boundaries, while thin lines serve only to illustrate the relative sizes of the grains.
III.1 Case $n=0$ ($l=4a$)
The present case is the simplest and is illustrated in Fig. 1. The initial prism contains $2^{0+2}=4$ unit cubes and has volume
$V_{\,0}^{\,0}=4a^{3}.$
Only one fragmentation stage is possible. The prism $T_{\,0}^{\,0}$ is divided into four equal parts and reassembled into a cross-shaped tower $T_{\,0}^{\,1}$ with a cavity at its centre, as shown in Fig. 1. The resulting volume is
$V_{\,0}^{\,1}=5a^{3},\qquad R_{\,0}^{\,1}=\frac{5}{4}.$
Thus, the maximal-volume configuration obtained by reassembling the fragments occupies a volume larger than that of the original object by a factor $5/4$.
The four cubes can also be rearranged into a compact configuration with no gaps, reproducing the original volume. The significance of this example is that random or unconstrained reassembly of the fragments can lead to a volume larger than that of the original object, while remaining bounded above by the maximal value $5/4$.
[…]
Figure 1: Fragmentation stages for case n=0(l=4a)n=0\,(l=4a). The initial prism $T_{\,0}^{\,0}$ and the resulting tower $T_{\,0}^{\,1}$ are shown in a cross-shaped arrangement.
In the following, the findings are summarised and rewritten in a form that is convenient for deriving general formulae for the volumes, relative volumes, and the total number of fragmentation stages.
[…]
$V_{\,0}^{\,1}$ $=4a^{3}+a^{3}=4\times 2^{0}a^{3}+2^{2\times 0}a^{3},$
$R_{\,0}^{\,1}$ $=\frac{5}{4},$
$T(0)$ $=0+1.$
$T(n)$ denotes the total number of fragmentation stages for an initial object indexed by $n$.
III.2 Case $n=1$ ($l=8a$)
In the case $n=1$ (see Fig. 2), the initial prism $T_{\,1}^{\,0}$ contains $2^{1+2}=8$ unit cubes and has volume
$V_{\,1}^{\,0}=8a^{3}.$
After the first fragmentation, which produces four pieces of length $2^{1}a=2a$, the maximal-volume tower $T_{\,1}^{\,1}$ has
$V_{\,1}^{\,1}=8a^{3}+4a^{3}=12a^{3},\qquad R_{\,1}^{\,1}=\frac{3}{2}.$
The quantity $4a^{3}$ corresponds to the volume of the cavity created at this stage.
Further division of each building block into two equal parts leads to the final stage, at which each block is a cube. These cubes are assembled into a two-level tower $T_{\,1}^{\,2}$ with a smaller cavity of volume $2a^{3}$,
$V_{\,1}^{\,2}=8a^{3}+2a^{3}=10a^{3},\qquad R_{\,1}^{\,2}=\frac{5}{4}.$
The resulting sequence of relative volumes is therefore
$R_{\,1}^{\,1}=\frac{3}{2}\;\longrightarrow\;R_{\,1}^{\,2}=\frac{5}{4}.$
[…]
Figure 2: Fragmentation stages for case n=1(l=8a)n=1\,(l=8a). The initial prism $T_{\,1}^{\,0}$ and the two resulting towers $T_{\,1}^{\,1}$ and $T_{\,1}^{\,2}$ are shown. The towers form a conjugate pair, with the second tower enclosing a smaller volume.
In the following, the findings are summarised in a form convenient for deriving general expressions for the volumes, relative volumes, and the total number of fragmentation stages.
$V_{\,1}^{\,0}$ $=8a^{3}=4\times 2^{1}a^{3},$ (2)
[…]
$V_{\,1}^{\,2}$ $=8a^{3}+2a^{3}=4\times 2^{1}a^{3}+2^{2\times 1-1}a^{3},$ (4)
$T(1)$ $=1+1.$ (5)
The tower $T_{\,1}^{\,1}$ is built from blocks of length $2a$, whereas the tower $T_{\,1}^{\,2}$ is constructed from unit blocks. These unit blocks are arranged such that the resulting structure appears as if it were built from the same type of building blocks as $T_{\,1}^{\,1}$.
This observation suggests that the tower $T_{\,1}^{\,2}$ may be regarded, at a coarse-grained geometric level, as being assembled from the same building blocks as $T_{\,1}^{\,1}$, differing only in their global arrangement. In this sense, one configuration may be transformed into the other by rearrangement of the building blocks, despite differences in their internal composition.
Situations of this type recur in all subsequent cases. It is therefore useful to introduce a specific terminology: a pair of towers that either appear to be, or are actually constructed from the same building blocks, but differ in their overall geometry and volume, are referred to as conjugate towers.
If no energetic preference is associated with either configuration, both may occur with comparable probability despite their different volumes. In this sense, the two configurations may be viewed as distinct “geometric phases,” one denser than the other.
III.3 Case $n=2$ ($l=16a$)
In the present case (see Fig. 3), the initial object $T_{\,2}^{\,0}$ consists of $16$ unit cubes arranged in a single row.
The first tower $T_{\,2}^{\,1}$ is obtained by dividing the initial object into four identical parts, each of length $4a$, and arranging the resulting building blocks as shown in Fig. 3. In this configuration, a cavity is formed at the centre of the tower. The total volume of the tower is
$V_{\,2}^{\,1}=32a^{3},$
where the volume of the cavity is $16a^{3}$.
By subsequently dividing each building block into two equal parts at each stage, two additional towers, $T_{\,2}^{\,2}$ and $T_{\,2}^{\,3}$, are obtained. Their corresponding volumes are
$V_{\,2}^{\,2}=24a^{3}\qquad V_{\,2}^{\,3}=20a^{3}.$
The resulting sequence of relative volumes is therefore
$2.00\to 1.50\to 1.25.$
[…]
It is observed that, relative to the initial object, the volume of the towers first increases at the initial fragmentation stage and then decreases monotonically at subsequent stages. This behaviour is governed by the cavity formed during the first rearrangement: while a cavity is created at the first stage, its volume decreases as the fragmentation proceeds.
A further parallel with the previous case may be drawn. The towers $T_{\,2}^{\,1}$ and $T_{\,2}^{\,3}$ may be regarded as being built from identical building elements of length $4a$, arranged in different geometric configurations. As before, such a pair is referred to as a pair of conjugate towers. The intermediate tower $T_{\,2}^{\,2}$ therefore lacks an obvious conjugate partner.
To clarify this point, the conjugation procedure is examined more closely. The tower $T_{\,2}^{\,3}$ can be obtained from $T_{\,2}^{\,1}$ by rotating the horizontally arranged building blocks into a vertical orientation. The same operation can be applied to the building blocks of the tower $T_{\,2}^{\,2}$: rotating the horizontally arranged blocks into a vertical position produces a tower that is identical in appearance to the original one, even though its internal construction differs. This behaviour is a consequence of the square shape of the tower’s lateral faces.
The tower $T_{\,2}^{\,2}$ is therefore referred to as a self-conjugate tower. It will be shown later that, whenever $n$ is even, the middle tower in the sequence is self-conjugate.
From this perspective, transitions between conjugate towers may be interpreted as transitions between distinct “geometric phases” with different densities. In contrast, the transition associated with a self-conjugate tower corresponds to a change between different symmetry realisations of the same geometric state. In this sense, transitions between conjugate towers resemble first-order–like phase transitions, while transitions involving self-conjugate towers may be viewed as second-order–like phase transitions.
In the case $n=3$, two pairs of towers are present, namely $(T_{\,3}^{\,1},T_{\,3}^{\,4})$ and $(T_{\,3}^{\,2},T_{\,3}^{\,3})$. Within each pair, a first-order-like geometrical phase transition is possible through a pure rearrangement of the building blocks. By contrast, for $n=4$ the tower pairs between which such geometrical phase transitions may occur are $(T_{\,4}^{\,1},T_{\,4}^{\,5})$ and $(T_{\,4}^{\,2},T_{\,4}^{\,4})$, while the tower $T_{\,4}^{\,3}$ is self-conjugate and is associated with a second-order-like phase transition.
These examples already reveal the general structure. For arbitrary $n$, the conjugate tower pairs are
$(T_{\,n}^{\,1},T_{\,n}^{\,n+1}),\,(T_{\,n}^{\,2},T_{\,n}^{\,n}),\,(T_{\,n}^{\,3},T_{\,n}^{\,n-1})\dots$
If $n$ is odd, there are $(n+1)/2$ such pairs, and every tower has a conjugate partner. If $n$ is even, there are $n/2$ conjugate pairs, while a single “middle” tower, $T_{n}^{n/2+1}$, remains self-conjugate. This odd – even distinction plays a central role in the classification of geometrical phase transitions and underlies the general formulae derived below.
It is now shown that the lateral faces of a self-conjugate tower are necessarily square. This requires the width and height of a lateral face to be equal. Since the width of a tower face is equal to the length of its building blocks, it is sufficient to demonstrate that this length coincides with the height of the tower.
Proposition 1.
The lateral faces of a self-conjugate tower are square.
Proof.
The equality of the width and height of a lateral face is established as follows.
First, a self-conjugate tower occurs only when $n$ is even and appears at the fragmentation stage $i=n/2+1$.
Second, the length of the building blocks at stage $i$ follows from the fragmentation rule. The initial length is divided by four at the first stage and is subsequently halved at each stage. The block length at stage $i$ is therefore
$l_{i}=\frac{2^{\,n+2}a}{2^{\,i+1}}=2^{\,n-i+1}a\,.$ (6)
Substitution of $i=n/2+1$ yields $l_{i}=2^{\,n/2}a$.
Third, the height of the tower is determined by the number of levels multiplied by $a$. A single level is formed at the first stage, and the number of levels doubles at each subsequent stage. The number of levels at stage $i$ is thus $2^{\,i-1}$, therefore height of the tower $h_{i}$ at stage $i$ is
$h_{i}=2^{\,i-1}a\,.$ (7)
Substitution of $i=n/2+1$ gives a tower height of $2^{\,n/2}a$.
Hence, the width and height of the lateral faces are therefore equal, and the lateral faces of a self-conjugate tower are square. ∎
Corollary 1.
The number of towers $T(n)$ created from an initial parental object with dimensions $2^{\,n+2}a\times a\times a$ is given by
T(n)=n+1.T(n)=n+1\,. (8)
Proof.
At the final stage of fragmentation, denoted by $f$, the block length $l_{f}$ is equal to $a$. Equation (6) therefore takes the form
$a=2^{\,n-f+1}a\,.$
It then follows directly that the final stage index is $f=n+1$. Consequently, the total number of towers produced during the fragmentation process is equal to $n+1$, which completes the proof. ∎
For each $n$, the sequence of towers $T_{\,n}^{\,i}$ exhibits a natural pairing: $T_{\,n}^{\,1}$ with $T_{\,n}^{\,n+1}$, $T_{\,n}^{\,2}$ with $T_{\,n}^{\,n}$, and so on. Towers within each pair are constructed from identical looking building blocks, but differ by a global rearrangement that leads to different enclosed cavity volumes.
Corollary 2.
The conjugate tower pairs are given by $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$, where $i=1,2,\dots,n+1$. That is, a tower at fragmentation stage $i$ is paired with the tower at fragmentation stage ni+2n-i+2.
Proof.
If a tower at fragmentation stage $i$ is paired with a tower at fragmentation stage $p$ (without loss of generality, $i\leqslant p$ may be assumed), then the length $l_{i}$ of the building blocks at stage $i$ is equal to the height $h_{p}$ of the tower at the paired stage $p$. Using Eqs. (6) and (7), this condition can be written as
$2^{\,n-i+1}a=2^{\,p-1}a\,.$
From this relation, the paired fragmentation stage is obtained as p=ni+2p=n-i+2. Consequently, the conjugate tower pairs are $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$. ∎
Corollary 3.
The number of conjugate tower pairs generated from an initial object indexed by $n$ depends on the parity of $n$. For odd $n$, the $n+1$ towers form $(n+1)/2$ conjugate pairs. For even $n$, the towers form $n/2$ conjugate pairs, and the remaining tower $T_{\,n}^{\,n/2+1}$ is self-conjugate.
Remark.
The emergence of conjugate and self-conjugate towers is a purely geometric consequence of the fragmentation – reassembly procedure. The term “geometric phase” is used here in an analogical sense, to denote distinct reassembled configurations of identical building blocks that enclose different volumes. No energetic or thermodynamic interpretation is implied.
III.4 General volume formula
Motivated by the explicit examples presented in the preceding subsections, a general expression for the maximal volume attained at each fragmentation stage is derived.
The volume of a tower is given by the sum of the material volume $2^{n+2}a^{3}$ of the initial parental object, which remains unchanged throughout the fragmentation process, and the volume of the enclosed space (cavity). The cavity volume is given by the product of the cross-sectional area of the cavity and the height of the tower, with the latter given by Eq. (7). The cross-sectional area of the cavity is equal to the square of the building-block length $l_{i}$, given by Eq. (6). The enclosed cavity volume is therefore expressed as
$l_{i}^{2}h_{i}=2^{\,2n-(i-1)}a^{3}.$
Proposition 2.
The volume $V_{\,n}^{\,i}$ of the tower $T_{\,n}^{\,i}$ reassembled at fragmentation stage $i$ from an initial parental object of dimensions $2^{\,n+2}a\times a\times a$ is given by
$V_{\,n}^{\,i}=2^{\,n+2}a^{3}+2^{\,2n-(i-1)}a^{3},\qquad i=1,2,\dots,n+1.$ (9)
The corresponding relative volume is therefore obtained as
$R_{\,n}^{\,i}=\dfrac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}=1+2^{\,n-i-1}.$ (10)
The variation of the relative volume as a function of the fragmentation stage $i$ for several values of $n$ is illustrated in Fig. 4. As established above, the final fragmentation occurs at stage $i=n+1$. The relative volume at the final stage is therefore universal and is given by
$R_{\,n}^{\,n+1}=1+2^{-2}=\frac{5}{4}.$ (11)
The value $5/4$ thus represents a universal upper bound for the relative volume obtained after complete fragmentation of any sufficiently long object with square cross section.
The plots further indicate that the first fragmentation stage produces the largest increase in volume, after which the relative volume decreases monotonically with increasing stage index. This behaviour is governed by the evolution of the central cavity: the initial object encloses no void space, whereas the first fragmentation generates the largest possible cavity. At each subsequent stage, the cross-sectional area of the cavity is reduced by a factor of four, while the height of the tower increases only by a factor of two. Consequently, the enclosed cavity volume is halved at each stage, while the relative volume decreases monotonically but is not halved because of the additive constant.
[…]
Figure 4: Relative volume $R_{\,n}^{\,i}$ as a function of fragmentation stage $i$ for several values of $n$. Each curve corresponds to a fixed initial length $l=2^{n+2}a$. The first fragmentation stage yields the maximal relative volume, followed by a monotonic decrease towards the universal limit $R=5/4$ at $i=n+1$.
III.5 Remarks on paired configurations and “phase” interpretation
As demonstrated below, the pairing structure is closely related to the symmetry properties of the tower configurations.
The relative-volume difference for a conjugate pair (i,n+2i)(i,n+2-i) is given by
$\Delta R_{\,n}^{\,i}$ $=R_{\,n}^{\,i}-R_{\,n}^{\,n+2-i}=\bigl(1+2^{\,n-i-1}\bigr)-\bigl(1+2^{\,n-(n+2-i)-1}\bigr)$
$=2^{\,n-i-1}-2^{\,i-3}.$ (12)
This difference is maximal for $i=1$ (corresponding to the pairing of the first and last towers) and decreases as the index $i$ approaches the central value, $i=n/2+1$.
The relative-volume difference vanishes when
$i=\frac{n}{2}+1\,.$ (13)
Since $i$ is a non-negative whole number, this condition requires careful interpretation. If $n$ is even, Eq. (13) yields an integer value, and the pair (i,n+2i)(i,n+2-i) reduces to $(n/2+1,n/2+1)$. In this case, two conjugate towers coincide and form a single self-conjugate tower. If $n$ is odd, such a reduction does not occur, and a self-conjugate tower never exists.
Conjugate towers are characterized by a rearrangement of identical building blocks. When these towers coincide in the even-$n$ case, the resulting self-conjugate tower inherits properties of both configurations. It can therefore be inferred that the self-conjugate tower may be constructed from the same building blocks in two distinct orientations, implying the presence of an internal symmetry, in agreement with the behaviour already observed in the example $n=2$ discussed in the subsection Case $n=2(l=4a)$.
These paired configurations provide a simple geometric analogue of distinct phases. While they are composed of identical building blocks, they differ by a global rearrangement that modifies the enclosed cavity volume. Rearrangement-induced transitions between conjugate towers thus correspond to transitions between geometrically distinct configurations with different effective densities.
The geometric transitions observed in the present model admit a natural interpretation in the language of phase transitions. The relative volume $R_{n}^{\,i}$ plays the role of a macroscopic order parameter distinguishing tower configurations. Transitions between conjugate towers involve discontinuous changes in $R_{n}^{\,i}$ and may therefore be regarded as first-order-like. In contrast, when $n$ is even, the middle tower is self-conjugate: rearrangement of the building blocks alters their internal orientation without changing the overall geometry or volume. This transition is associated with a discrete rotational symmetry and corresponds to a change in internal symmetry rather than in macroscopic structure. In this sense, the self-conjugate tower represents a geometric analogue of a second-order phase transition.
In summary, the fragmented tower configurations are organized into conjugate pairs distinguished by a discrete symmetry of relative volume, with a self-conjugate configuration arising only for even $n$. This structure provides a geometric analogue of phase behavior, in which conjugate towers represent distinct macroscopic phases, while the self-conjugate tower corresponds to a symmetry-driven transition without a change in density.
IV Physical interpretation and connections to granular matter
To connect the present model with real granular systems, Eq. (10) is employed to evaluate the relative volume accessible to grains in experiment, while Eq. (12) is employed to identify the geometric phase transitions predicted by the model. Both expressions are formulated in terms of the discrete parameters $n$ and $i$, which must therefore be re-expressed in terms of experimentally measurable grain characteristics.
Throughout this section, it is assumed that grains do not interact with one another, that frictional effects are negligible, and that individual grains may be regarded as effectively weightless, such that no load is transmitted through the assembly. These assumptions are adopted in order to isolate the purely geometric contributions to volume evolution that are central to the present model.
IV.1 Remarks on experimental observation of the relative volume
A granular sample composed of identical grains with square cross-section of side length $a$ is considered. It is further assumed that these grains may be regarded as having been produced by sequential slicing of a hypothetical elongated parent object, in accordance with the fragmentation rule introduced earlier. By comparing the total mass of the granular sample with the mass of an individual grain, the total number of grains in the sample, denoted by $N_{g}$, is determined. Let $l_{g}$ denote the length of an individual grain. The length $l$ of the hypothetical parent object is then given by
$l=N_{g}\,l_{g}\,.$ (14)
Within the model, the same length is expressed by Eq. (1) in terms of the discrete parameter $n$. By equating the two expressions and solving for $n$, one obtains
$n=\log_{2}\!\left(\frac{N_{g}\,l_{g}}{a}\right)-2\,.$ (15)
Here, the parameter $a$ is assigned the same meaning in both the model description and the experimental system.
Our next goal is to identify the fragmentation stage of the hypothetical parental object at which the size of a model grain coincides with that of the grains used in the experiment. At fragmentation stage $i$, the grain length $l_{i}$ (given by Eq. (6)) must match the experimental grain length $l_{g}$, so that
$2^{\,n-i+1}a=l_{g}\,.$
Using Eq. (15), this condition yields
$i=\log_{2}N_{g}-1\,.$ (16)
As a consistency check, it is noted that for a single grain ($N_{g}=1$), one obtains $i=0$, corresponding to the absence of fragmentation. Substituting Eqs. (15) and (16) into Eq. (10) yields
$R=1+\frac{l_{g}}{4a}\,.$ (17)
This result demonstrates that the relative volume is determined exclusively by the geometric parameters of the grains and is independent of the details of the fragmentation history. All dependence on the hypothetical parent object is eliminated, rendering the indices $n$ and $i$ redundant; they are therefore omitted from Eq. (17).
Although the relative volume $R$ is expressed in Eq. (17) as a single function of the geometric ratio $l_{g}/a$, this quantity should be understood as a characteristic upper bound rather than as a uniquely reproducible experimental outcome. Under nominally identical preparation conditions, repeated realizations of the system are expected to yield a distribution of measured relative volumes clustered below this bound. For a granular assembly composed of grains with square cross section, the relative volume may therefore be determined experimentally by measuring only the grain length $l_{g}$ and the side length $a$ of the cross section. Upon successive fragmentation, these measurements may be repeated until the terminal stage is reached, at which $l_{g}=a$. In this limit, the relative volume approaches the value $5/4$, in agreement with Eq. (11). Exact quantitative agreement with the theoretical curve is not expected under experimental conditions. Nevertheless, the measured values are expected to follow the predicted qualitative trend and to remain bounded above by the maximal-volume curve provided by the model. Assuming square cross sections, these considerations may be summarised as follows.
Theorem (Liza’s theorem).
If a grain has length $l_{g}$ and a square cross section with side length $a$, the upper bound of the relative volume is determined solely by the ratio $l_{g}/a$.
Remark (Liza’s limit).
The upper bound of the relative volume decreases monotonically as the grain length $l_{g}$ decreases. Its absolute minimum is attained when $l_{g}=a$, where the upper bound equals $5/4$, a universal value referred to as Liza’s limit or Liza’s number. The upper bound always exceeds unity, corresponding to the volume obtained by packing all grains adjacently without voids.
Equation (17) may be tested against existing experimental data. In Ref. [9], experiments were conducted on monodisperse nylon $6/6$ cylindrical rods with diameter $1.8\,\mathrm{mm}$ and length $7.0\,\mathrm{mm}$. Approximately $7200$ rods were poured into a vertical glass tube of height $1\,\mathrm{m}$ and inner diameter $1.90\,\mathrm{cm}$. Although the present model is formulated for grains with square cross section, a reasonable estimate is expected to be obtained for cylindrical rods of comparable transverse dimensions. The rods are therefore approximated as grains with square cross section of side length a=1.8a=1.8mm and grain length $l_{g}=7.0\,\mathrm{mm}$. Substitution into Eq. (17) then yields
$R=1+\frac{l_{g}}{4a}\simeq 1.97\,.$
Experimentally, the quantity reported is the packing factor $\rho$ (also commonly denoted $\phi$), defined as the fraction of the available volume occupied by the particles. When the longest grain dimension is small compared to the characteristic dimensions of the container — in the present case, the rod length relative to the tube diameter — the packing factor may be approximated as $\rho\simeq R^{-1}$. This estimate gives $\rho\simeq 0.51$, which lies within the experimentally observed range $\rho=0.49$$0.55$ reported in Ref. [9].
IV.2 Remarks on experimental observation of the “phase transition”
An important question remains: can the geometric, rearrangement-induced phase transitions described above be observed experimentally? In addressing this question, two physical constraints must be taken into account. First, once a grain has been sliced, cut, or crushed, the resulting fragments cannot spontaneously reattach to reconstitute the original solid grain. Second, the levels of a tower are not mechanically bonded to one another; instead, the upper levels merely rest on the lower ones.
A tower at fragmentation stage $i$ is constructed from grains of length $l_{i}=2^{\,n-i+1}a$ and has a height $h_{i}=2^{\,i-1}a$ (see Eqs. (6) and (7)). The conjugate tower at stage ni+2n-i+2 is instead constructed from grains of length $l_{n-i+2}=2^{\,i-1}a$ and has a height $h_{n-i+2}=2^{\,n-i+1}a$. At first sight, the two towers appear to be assembled from identical grains: in the former case the grains are arranged horizontally, whereas in the latter they are aligned vertically. Crucially, however, this apparent equivalence is misleading. The vertically oriented “grain” is not a single object, but rather a stack of smaller fragments, whereas the horizontally oriented grain is an intact solid piece.
For illustration, the case $n=2$, shown in Fig. 3, may be considered. At fragmentation stage $i=1$, the grain length is $l_{1}=2^{\,2-1+1}a=4a$, and the tower consists of a single level of height $h_{1}=2^{\,1-1}a=a$. At the conjugate stage ni+2=3n-i+2=3, the grain length is $l_{3}=2^{\,i-1}a=a$, and the tower height is $h_{3}=2^{\,2-1+1}a=4a$. Four unit blocks are stacked vertically, producing an object that appears identical, at a coarse-grained geometric level, to a single grain of length $4a$ at stage i=1i=1\,.
As discussed earlier, a tower at fragmentation stage $i$ is paired with a tower at stage ni+2n-i+2\,. This pairing reflects the fact that the two configurations are composed of building blocks that are geometrically indistinguishable at the level of their external shape, despite differing in their internal structure. Without loss of generality, the case i<ni+2i<n-i+2 is considered. In this situation, the grains at stage ni+2n-i+2 are smaller than those at stage $i$. Although their vertical alignment reproduces the geometry of the larger grains, recombination into the intact grains of stage $i$ is not possible. By contrast, the grains at stage $i$ can be rearranged to form the tower corresponding to stage ni+2n-i+2\,.
It is therefore concluded that the geometric phase transition is intrinsically directional and may occur only in the direction
$i\;\longrightarrow\;n-i+2\,$,
unless the grains at stage $i$ are already arranged and oriented in the geometric configuration corresponding to stage ni+2n-i+2\,.
With this constraint in mind, the following experimental situation may be considered. Suppose that $N_{g}$ grains, each of length $l_{g}$, are given. The fragmentation stage $i$ of the hypothetical parental object corresponding to this configuration is determined by Eq. (16). The conjugate stage $i_{c}\equiv n-i+2\,$, expressed directly in terms of experimentally measurable parameters, is then given by
$i_{c}=\log_{2}\!\left(\frac{l_{g}}{a}\right)+1\,.$ (18)
If
$i<i_{c}\,$, (19)
the occurrence of a geometric phase transition is permitted; otherwise, such a transition is prohibited.
By substituting the expressions for $n$ and $i$ from Eqs. (15) and (16) into Eq. (12) for the difference between the relative volumes of conjugate configurations, the following expression is obtained:
$\Delta R=\frac{l_{g}}{4a}-\frac{1}{16}N_{g}\,.$ (20)
Here, the indices on $\Delta R$ are redundant and have therefore been omitted.
As already assumed above, the comparison is made under the convention i<ni+2i<n-i+2\,, adopted without loss of generality. Since the relative volume decreases monotonically with increasing fragmentation stage, the difference $\Delta R\,$ is therefore non-negative. Equation (20) then yields the criterion
$\frac{N_{g}a}{4\,l_{g}}\leqslant 1\,.$ (21)
Equality corresponds to the self-conjugate configuration, in which the transition involves a change in internal geometric symmetry without any change in the enclosed volume.
If condition (21) is not satisfied, the number of grains $N_{g}$ may, in principle, be reduced (provided $N_{g}\gg 1$) until the criterion is met. Combining these requirements yields
$1\ll N_{g}\leqslant\frac{4l_{g}}{a}\,$,
indicating that the occurrence of a geometric phase transition requires grains with a sufficiently large aspect ratio.
The implications of this criterion may be illustrated by reconsidering the example of the nylon rods discussed above. In this case, the phase-transition condition implies $N_{g}<16\,$. As a further example with a larger aspect ratio, matchsticks (with tips removed), characterised by a cross-sectional side length $a=2\,\mathrm{mm}$ and a length $l_{g}=5\,\mathrm{cm}\,$, yield the condition $N_{g}<100\,$. In both cases, the number of grains required for the observation of the transition is relatively small.
It is therefore suggested that such geometric phase transitions are restricted to mesoscopic systems, which may explain why they are not commonly observed in everyday granular materials.
The phase-transition criterion (21) also provides additional insight into the interpretation of Eq. (17). When this criterion is not satisfied, the system admits a unique geometric configuration, and repeated experimental measurements of the relative volume $R$ are expected to form a single cluster centered around the characteristic value prescribed by Eq. (17). If the criterion is satisfied, however, two conjugate geometric configurations become accessible. In this regime, repeated preparations of the system may populate either configuration, resulting in two distinct clusters in the measured values of $R$, each corresponding to one of the conjugate geometric phases.
The observations reported in Ref. [4] support the notion that geometric phase transitions of the type discussed here may be induced by mechanical agitation. In that study, pistachio shells - whose shapes deviate significantly from the idealized grains considered in the present model - were found to interlock, nest, and compact under shaking. This behaviour indicates that such transitions are unlikely to occur uniformly throughout the sample, but instead arise locally within small clusters of grains. This naturally motivates consideration of clusters whose grain number and aspect ratio satisfy the phase-transition criterion derived above. Local rearrangements of this type may then cumulatively contribute to the macroscopic volume reduction observed during agitation. To test this hypothesis experimentally, the bulk packing density may be modelled as a superposition of contributions from different geometric configurations (or phases), with the evolution of this distribution monitored as a function of the agitation protocol.
Although the present model is purely geometric and entirely static in nature, any experimental realization necessarily involves real grains subject to friction, gravity, and contact constraints. Gentle mechanical agitation is therefore introduced not to endow the system with thermodynamic character, but simply to facilitate internal rearrangements and allow the system to explore its admissible geometric configurations. In this sense, agitation serves as a practical means of reducing kinetic trapping and history dependence, rather than as a source of thermalization or energetic driving.
Experimental protocol (conceptual).
A minimal experimental test of the proposed geometric phase transition could proceed as follows. A collection of $N_{g}\gg 1$ elongated grains of uniform length $l_{g}$ and square cross section $a\times a$ (e.g., rods or trimmed matchsticks) is prepared. The grains are assembled in a loosely confined container that permits internal rearrangement under external agitation while suppressing large-scale translation of the sample as a whole. For fixed values of $N_{g}$ and $l_{g}$, the system is repeatedly prepared under identical macroscopic conditions and subjected to controlled mechanical excitation (tapping or shaking) of moderate intensity. Such excitation is intended solely to overcome frictional constraints and enable rearrangements between geometric configurations, rather than to induce thermal equilibration. After each preparation, the bulk occupied volume is measured and the corresponding relative volume $R$ is computed.
If the criterion (21) is not satisfied, the measurements are expected to yield a single reproducible cluster of $R$ values. If the criterion is satisfied, two distinct clusters of $R$ values may be observed across repeated preparations, corresponding to the paired conjugate configurations. The appearance of such bi-modality would indicate a geometric phase transition driven by configurational rearrangement rather than by changes in material properties. The separation between the two clusters should correlate with the prediction of Eq. (20).
These considerations should be taken into account when testing the dependence of the relative volume on grain size, Eq. (17). In the parameter regime where the criterion (21) is satisfied, two distinct branches are expected, corresponding to conjugate configurations. Outside this regime, where the criterion is not satisfied, a single branch should be observed.
V Summary
A geometric model of granular matter composed of grains with square cross section has been developed by representing the system as originating from a hypothetical elongated parent object. The characteristic size of this object is fixed by the number and dimensions of the constituent grains.
Fragmentation is implemented through a well-defined slicing algorithm, applied iteratively and described in detail in the main text. At each stage of fragmentation, the resulting fragments are reassembled into configurations of maximal admissible volume, referred to as towers. The volumes of these towers were analysed as functions of the fragmentation stage. In realistic granular systems composed of elongated grains, the constituents at a given stage of fragmentation are expected to arrange in a largely unconstrained manner, resulting in occupied volumes that lie between those of compact packing and the largest geometrically admissible configurations. The construction of maximal-volume towers therefore provides a purely geometric upper bound and allows the characteristic scale and qualitative trends of volume changes induced by fragmentation to be estimated.
It has been shown that, at the initial stage of fragmentation, the occupied volume increases relative to that of the parent object. With further fragmentation, the volume then decreases monotonically, approaching at the final stage the universal value $5/4$ of the original volume. It follows that, throughout the fragmentation process, a material composed of elongated grains typically occupies a volume larger than that of the original object, unless the grains are packed perfectly without voids. Within the fragmentation sequence, pairs of configurations have been identified and termed conjugate towers. These towers are composed of geometrically indistinguishable building blocks but enclose different volumes. When the number of fragmentation stages is even, a self-conjugate tower appears at the centre of the sequence. The magnitude of the volume difference between conjugate towers depends on their separation along the fragmentation sequence: towers corresponding to nearby stages exhibit smaller differences, whereas the largest difference occurs between the first and final stages of fragmentation.
Explicit relations have also been derived, and a procedure has been outlined for connecting the idealised geometric model to systems of real cuboidal grains and for experimentally testing its predictions. These include the monotonic evolution of volume with fragmentation and the emergence of phase-like transitions between distinct packing configurations driven purely by geometric rearrangement.
In summary, an analytically tractable geometric model of fragmentation and reassembly has been presented, revealing universal features of volume evolution and phase-like behaviour in granular systems. Despite its simplicity, the model captures essential geometric aspects of granular matter and provides a useful framework for the investigation of packing geometry, porosity, and metastable configurations.
Acknowledgements
This work is dedicated to the memory of the author’s grandmother, Liza. Many years ago, she remarked that a given mass of corn kernels occupies a smaller volume than the flour obtained from them. This simple observation provided the original motivation for the present study and inspired the investigation of the geometric origins underlying such volume changes.
The author thanks T. Börzsönyi for valuable comments on the role of Van der Waals forces in fine powder packing and for helpful discussions during the preparation of this manuscript.