### v2

# Investigating Disordered Granular Matter via Ordered Geometric Fragmentation

Abstract
The evolution of occupied volume under progressive fragmentation of granular matter is studied using a purely geometric model. Rather than modelling disorder directly, properties are investigated by analysing highly ordered reference configurations that provide sharp upper bounds on accessible volume. Grains are idealised as fragments from a hypothetical elongated parent prism with square cross section, sequentially sliced and reassembled into configurations that maximise enclosed volume.
Analytic expressions are derived for the maximal volume at each fragmentation stage. Volume evolution is non-monotonic: initial fragmentation produces structures exceeding the original volume, while further fragmentation leads to monotonic decrease converging to 5/4 times the initial volume, independent of fragment number. The packing fraction obeys the asymptotic scaling law of inverse proportionality to aspect ratio, in agreement with experimental observations.
The model reveals pairs of configurations built from geometrically indistinguishable building blocks yet enclosing different volumes. These conjugate configurations constitute geometric analogues of distinct phases connected by rearrangement-induced transitions. A criterion for observability is derived, showing such transitions are restricted to systems of limited grain number but may occur locally as domain formation in larger assemblies. Comparison with experimental data confirms the model provides a lower bound on packing fraction and predicts domain sizes should scale linearly with aspect ratio, testable through X-ray tomography.
I Introduction
Packing problems arise in a wide range of physical contexts, including granular compaction, powder processing, jamming phenomena, and colloidal self-assembly. A substantial body of work has therefore been devoted to understanding how particle shape, anisotropy, and geometric constraints influence packing density and available volume; see, for example, Refs. [1, 2, 3] and references therein. In most studies, the primary objective is to determine optimal or typical packing fractions for prescribed particle shapes within disordered or statistically defined ensembles.
Recent theoretical progress has been made in understanding random close packing (RCP) of monodisperse spherical particles through analytical frameworks linking packing fraction to coordination number and the onset of mechanical rigidity [4]. These approaches identify the jamming transition as a critical point characterized by a threshold coordination number, with packing fractions that are largely independent of system size. However, the case of non-spherical elongated particles remains analytically challenging and has been explicitly identified as an open problem requiring further investigation [4]. The present work contributes to this area through a purely geometric approach that reveals how fragmentation-induced volume evolution in elongated grain assemblies can be understood from geometric constraints alone, without recourse to mechanical or energetic considerations.
In contrast to conventional packing studies, the present work addresses a complementary question: how does the volume occupied by granular matter evolve under progressive fragmentation? Rather than modelling disorder directly, this question is examined through a controlled geometric construction based on a single elongated parent object that is sequentially partitioned into identical building blocks. At each fragmentation stage, the maximal enclosed volume attainable after reassembly is evaluated. The resulting configurations are not intended to represent typical granular packings; instead, they provide sharp geometric bounds that elucidate systematic trends in volume evolution induced by fragmentation.
[…]
The packing of elongated particles has been studied extensively both experimentally and theoretically [7, 8, 9, 10]. Systematic experiments have demonstrated that packing fraction decreases with increasing aspect ratio for randomly packed rods [9, 11], with values ranging from $\phi\approx 0.55$ for short rods (aspect ratio $\alpha\approx 4$) to $\phi\approx 0.22$ for highly elongated particles ($\alpha\approx 32$). Importantly, experimental measurements reveal that nominally identical particles under similar preparation conditions can yield a range of packing fractions rather than a single reproducible value [10], indicating the presence of multiple metastable configurations — a phenomenon that finds a natural geometric interpretation in the present model.
While much attention has been devoted to the packing of particles with fixed shapes, the influence of fragmentation on packing density remains less systematically explored. Fragmentation processes — whether mechanical comminution, crushing, or controlled slicing — alter both the size distribution and the geometric properties of constituent grains [12]. The resulting changes in particle aspect ratio and surface morphology can significantly modify packing efficiency. The present work isolates the purely geometric contribution to these effects by considering idealized fragmentation of a single parent object into progressively smaller, identical building blocks.
Building on these ideas, a minimal, fully analytic model is constructed that isolates the geometric consequences of fragmentation and reassembly. Explicit expressions for the maximal enclosed volume at each fragmentation stage are obtained, and several structural features are revealed. In particular, the emergence of paired (conjugate) configurations composed of geometrically indistinguishable building blocks, yet enclosing different volumes, is identified. The model predicts bounds on packing density that are shown to be consistent with experimental scaling laws for elongated particles. Furthermore, the geometric framework reveals local ordering phenomena in bulk granular assemblies that can be tested through X-ray tomography experiments.
The objectives of this paper are as follows: (i) to quantify the evolution of volume under sequential fragmentation of granular matter using a simple geometric model; (ii) to investigate structural relations among configurations reassembled from differently fragmented building blocks; (iii) to derive explicit formulae that connect the parameters of the model to experimentally measurable quantities; and (iv) to compare theoretical predictions with experimental data and propose testable signatures of the geometric mechanisms identified here.
II Model and notation
In this work, the geometric properties of granular matter composed of idealised grains are investigated within a purely static geometric framework. Friction between grains is assumed to be negligible, and the weight of individual grains is ignored. All grains are assumed to originate from a single, very long hypothetical parent object. The number of grains produced and their individual sizes are determined entirely by the geometry of this initial object.
Grains are modelled as rectangular parallelepipeds with square cross-section. Accordingly, a family of long rectangular prisms (hereafter referred to as initial objects) is considered, denoted by $T_{\,n}^{\,0}$, with square cross-section $a\times a$ and length
$l=2^{\,n+2}a,\qquad n=0,1,2,\dots.$ (1)
(The length is written in this particular form for convenience in subsequent calculations.)
Each initial object may conveniently be viewed as consisting of
$N_{\rm tot}=2^{\,n+2}$ (2)
unit cubes of side length $a$. The corresponding volume is
$V_{\,n}^{\,0}=2^{\,n+2}a^{3}.$ (3)
Fragmentation of the initial object is carried out stage by stage. At each stage $i=1,2,\dots,n+1$, the object is fragmented according to a rule defined below, and the resulting pieces (grains) are reassembled into a configuration that maximises the volume. The reassembled object is referred to as a tower and is denoted by $T_{\,n}^{\,i}$ , with volume $V_{\,n}^{\,i}$ and relative volume
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}.$ (4)
In this notation, the lower index $n$ characterises the size of the initial parent object via Eq. (1), while the upper index $i$ labels the fragmentation stage.
II.1 Fragmentation rule
The fragmentation rule is defined recursively as follows. Stage 0 denotes the initial configuration prior to any fragmentation, while stages $i=1,2,\dots,n+1$ correspond to successive fragmentation steps.
  1. 1.
    Stage 0: The initial prism $T_{\,n}^{\,0}$ is defined as a rectangular prism of dimensions $2^{\,n+2}a\times a\times a$.
  2. 2.
    Stage 1: The prism is divided into four equal parts of length $l/4=2^{\,n}a$. These parts are reassembled to maximise the enclosed volume, yielding the tower $T_{\,n}^{\,1}$.
  3. 3.
    Stage 2: Each piece obtained at Stage 1 is divided along its longest axis into two equal parts. The resulting eight pieces are reassembled into the tower $T_{\,n}^{\,2}$.
  4. 4.
    This procedure is continued by dividing each piece along its longest axis into two equal parts at each subsequent stage, until stage $n+1$, at which point all pieces are unit cubes and the maximal-volume configuration $T_{\,n}^{\,n+1}$ is obtained.
After fragmentation, the elementary building blocks can, in general, be assembled into towers whose cross sections form parallelograms. In the present model, the cross section is restricted to be square. This choice is motivated by the fact that, among all parallelograms with a given perimeter, the square encloses the maximal cross-sectional area.
The square geometry therefore represents an extremal, volume-maximising configuration within this class and serves as a natural reference case. No claim is made that this construction exhausts all possible configurations, nor that a global maximal-volume theorem is established.
III Fragmentation – reassembly process and volume evolution
Using the fragmentation rule introduced in Sec. II.1, the slicing of long objects is presented explicitly below for the first few values of $n$. For each $n$, the sequence of towers $T_{\,n}^{\,i}$ produced at fragmentation stages $i=1,\dots,n+1$, including the initial object $i=0$, is listed alongside their volumes $V_{\,n}^{\,i}$ and corresponding relative volumes $R_{\,n}^{\,i}$. These examples motivate the general formulae, which are subsequently derived using mathematical induction.
In the figures that follow, thick lines indicate grain boundaries, while thin lines serve only to illustrate the relative sizes of the grains.
III.1 Case $n=0$ ($l=4a$)
The present case is the simplest and is illustrated in Fig. 1. The initial prism contains $2^{0+2}=4$ unit cubes and has volume
$V_{\,0}^{\,0}=4a^{3}.$
Only one fragmentation stage is possible. The prism $T_{\,0}^{\,0}$ is divided into four equal parts and reassembled into a cross-shaped tower $T_{\,0}^{\,1}$ with a cavity at its centre, as shown in Fig. 1. The resulting volume is
$V_{\,0}^{\,1}=5a^{3},\qquad R_{\,0}^{\,1}=\frac{5}{4}.$
Thus, the maximal-volume configuration obtained by reassembling the fragments occupies a volume larger than that of the original object by a factor $5/4$.
The four cubes can also be rearranged into a compact configuration with no gaps, reproducing the original volume. The significance of this example is that random or unconstrained reassembly of the fragments can lead to a volume larger than that of the original object, while remaining bounded above by the maximal value $5/4$.
[…]
Figure 1: Fragmentation stages for case n=0(l=4a)n=0\,(l=4a). The initial prism $T_{\,0}^{\,0}$ and the resulting tower $T_{\,0}^{\,1}$ are shown in a cross-shaped arrangement.
In the following, the findings are summarised and rewritten in a form that is convenient for deriving general formulae for the volumes, relative volumes, and the total number of fragmentation stages. $T(n)$ denotes the total number of fragmentation stages for an initial object indexed by $n$, so that
[…]
$V_{\,0}^{\,1}$ $=4a^{3}+a^{3}=4\times 2^{0}a^{3}+2^{2\times 0}a^{3},$
$R_{\,0}^{\,1}$ $=\frac{5}{4},$
$T(0)$ $=0+1.$
III.2 Case $n=1$ ($l=8a$)
In the case $n=1$ (see Fig. 2), the initial prism $T_{\,1}^{\,0}$ contains $2^{1+2}=8$ unit cubes and has volume
$V_{\,1}^{\,0}=8a^{3}.$
After the first fragmentation, which produces four pieces of length $2^{1}a=2a$, the maximal-volume tower $T_{\,1}^{\,1}$ has
$V_{\,1}^{\,1}=8a^{3}+4a^{3}=12a^{3},\qquad R_{\,1}^{\,1}=\frac{3}{2}.$
The quantity $4a^{3}$ corresponds to the volume of the cavity created at this stage.
Further division of each building block into two equal parts leads to the final stage, at which each block is a cube. These cubes are assembled into a two-level tower $T_{\,1}^{\,2}$ with a smaller cavity of volume $2a^{3}$,
$V_{\,1}^{\,2}=8a^{3}+2a^{3}=10a^{3},\qquad R_{\,1}^{\,2}=\frac{5}{4}.$
The resulting sequence of relative volumes is therefore
$R_{\,1}^{\,1}=\frac{3}{2}\;\longrightarrow\;R_{\,1}^{\,2}=\frac{5}{4}.$
[…]
Figure 2: Fragmentation stages for case n=1(l=8a)n=1\,(l=8a). The initial prism $T_{\,1}^{\,0}$ and the two resulting towers $T_{\,1}^{\,1}$ and $T_{\,1}^{\,2}$ are shown. The towers form a conjugate pair, with the second tower enclosing a smaller volume.
In the following, the findings are summarised in a form convenient for deriving general expressions for the volumes, relative volumes, and the total number of fragmentation stages.
$V_{\,1}^{\,0}$ $=8a^{3}=4\times 2^{1}a^{3},$
[…]
$V_{\,1}^{\,2}$ $=8a^{3}+2a^{3}=4\times 2^{1}a^{3}+2^{2\times 1-1}a^{3},$
$T(1)$ $=1+1.$
The tower $T_{\,1}^{\,1}$ is built from blocks of length $2a$, whereas the tower $T_{\,1}^{\,2}$ is constructed from unit blocks arranged such that the resulting structure appears, at a coarse-grained geometric level, as if it were assembled from the same building blocks as $T_{\,1}^{\,1}$, differing only in their global arrangement. In this sense, one configuration may be transformed into the other by rearrangement of the building blocks, despite differences in their internal composition.
Situations of this type recur in all subsequent cases. It is therefore useful to introduce a specific terminology: a pair of towers that either appear to be, or are actually constructed from the same building blocks, but differ in their overall geometry and volume, are referred to as conjugate towers.
Within the purely geometric framework of the present model, if no energetic preference is associated with either configuration, both may occur with comparable probability despite their different volumes. In this sense, the two configurations may be viewed as distinct ‘geometric phases’, one denser than the other.
III.3 Case $n=2$ ($l=16a$)
In the present case (see Fig. 3), the initial object $T_{\,2}^{\,0}$ consists of $16$ unit cubes arranged in a single row.
The first tower $T_{\,2}^{\,1}$ is obtained by dividing the initial object into four identical parts, each of length $4a$, and arranging the resulting building blocks as shown in Fig. 3. In this configuration, a cavity is formed at the centre of the tower. The total volume of the tower is
$V_{\,2}^{\,1}=32a^{3},$
where the volume of the cavity is $16a^{3}$.
By subsequently dividing each building block into two equal parts at each stage, two additional towers, $T_{\,2}^{\,2}$ and $T_{\,2}^{\,3}$, are obtained. Their corresponding volumes are
$V_{\,2}^{\,2}=24a^{3},\qquad V_{\,2}^{\,3}=20a^{3}.$
The resulting sequence of relative volumes is therefore
$R_{\,2}^{\,1}=2\;\longrightarrow\;R_{\,2}^{\,2}=\frac{3}{2}\;\longrightarrow\;R_{\,2}^{\,3}=\frac{5}{4}.$
[…]
Relative to the initial object, the volume of the towers first increases at the initial fragmentation stage and then decreases monotonically at subsequent stages. This behaviour is governed by the cavity formed during the first rearrangement: while a cavity is created at the first stage, its volume decreases as the fragmentation proceeds.
A further parallel with the case $n=1$ (Sec. III.2) may be drawn. The towers $T_{\,2}^{\,1}$ and $T_{\,2}^{\,3}$ may be regarded as being built from identical building elements of length $4a$, arranged in different geometric configurations. As before, such a pair is referred to as a pair of conjugate towers. The intermediate tower $T_{\,2}^{\,2}$ does not possess a conjugate partner in the same sense.
To clarify this point, the conjugation procedure is examined more closely. The tower $T_{\,2}^{\,3}$ can be obtained from $T_{\,2}^{\,1}$ by rotating the horizontally arranged building blocks into a vertical orientation. The same operation can be applied to the building blocks of the tower $T_{\,2}^{\,2}$: rotating the horizontally arranged blocks into a vertical position produces a tower that is identical in appearance to the original one, even though its internal construction differs. This behaviour is a consequence of the square shape of the tower’s lateral faces.
The tower $T_{\,2}^{\,2}$ is therefore referred to as a self-conjugate tower. As shown in Proposition 1 below, whenever $n$ is even, the middle tower in the sequence is self-conjugate.
From this perspective, transitions between conjugate towers may be interpreted as transitions between distinct ‘geometric phases’ with different densities. In contrast, the transition associated with a self-conjugate tower corresponds to a change between different symmetry realisations of the same geometric state. In this sense, transitions between conjugate towers resemble first-order-like phase transitions, while transitions involving self-conjugate towers may be viewed as second-order-like phase transitions.
In the case $n=3$, two pairs of towers are present, namely $(T_{\,3}^{\,1},T_{\,3}^{\,4})$ and $(T_{\,3}^{\,2},T_{\,3}^{\,3})$. Within each pair, a first-order-like geometrical phase transition is possible through a pure rearrangement of the building blocks. By contrast, for $n=4$ the tower pairs between which such geometrical phase transitions may occur are $(T_{\,4}^{\,1},T_{\,4}^{\,5})$ and $(T_{\,4}^{\,2},T_{\,4}^{\,4})$, while the tower $T_{\,4}^{\,3}$ is self-conjugate and is associated with a second-order-like phase transition.
These examples already reveal the general structure. For arbitrary $n$, the conjugate tower pairs are
$(T_{\,n}^{\,1},T_{\,n}^{\,n+1}),\,(T_{\,n}^{\,2},T_{\,n}^{\,n}),\,(T_{\,n}^{\,3},T_{\,n}^{\,n-1}),\,\dots$ (5)
If $n$ is odd, there are $(n+1)/2$ such pairs, and every tower has a conjugate partner. If $n$ is even, there are $n/2$ conjugate pairs, while a single ‘middle’ tower, $T_{\,n}^{\,n/2+1}$, remains self-conjugate. This odd–even distinction plays a central role in the classification of geometrical phase transitions and underlies the general formulae derived in Sec. III.4.
The following proposition establishes that the lateral faces of a self-conjugate tower are necessarily square, as claimed above. This requires the width and height of a lateral face to be equal. Since the width of a tower face is equal to the length of its building blocks, it is sufficient to demonstrate that this length coincides with the height of the tower.
Proposition 1.
The lateral faces of a self-conjugate tower are square.
Proof.
The equality of the width and height of a lateral face is established as follows.
First, a self-conjugate tower occurs only when $n$ is even and appears at the fragmentation stage $i=n/2+1$.
Second, the length of the building blocks at stage $i$ follows from the fragmentation rule. The initial length is divided by four at the first stage and is subsequently halved at each stage. The block length at stage $i$ is therefore
$l_{i}=\frac{2^{\,n+2}a}{2^{\,i+1}}=2^{\,n-i+1}a\,.$ (6)
Substitution of $i=n/2+1$ into Eq. (6) yields $l_{\,n/2+1}=2^{\,n/2}a$.
Third, the height of the tower is determined by the number of levels multiplied by $a$. A single level is formed at the first stage, and the number of levels doubles at each subsequent stage. The number of levels at stage $i$ is thus $2^{\,i-1}$, therefore the height of the tower $h_{i}$ at stage $i$ is
$h_{i}=2^{\,i-1}a\,.$ (7)
Substitution of $i=n/2+1$ into Eq. (7) gives $h_{\,n/2+1}=2^{\,n/2}a$.
Hence, the width and height of the lateral faces are equal, and the lateral faces of a self-conjugate tower are square. ∎
Corollary 1.
The number of towers $T(n)$ created from an initial parent object with dimensions $2^{\,n+2}a\times a\times a$ is given by
T(n)=n+1.T(n)=n+1\,. (8)
Proof.
At the final stage of fragmentation, denoted by $f$, the block length $l_{f}$ is equal to $a$. Equation (6) therefore takes the form
$a=2^{\,n-f+1}a\,.$
It then follows directly that the final stage index is $f=n+1$. Since towers are produced at stages $i=1,2,\dots,f$, the total number of towers produced during the fragmentation process is equal to $n+1$, which completes the proof. ∎
For each $n$, the sequence of towers $T_{\,n}^{\,i}$ exhibits a natural pairing: $T_{\,n}^{\,1}$ with $T_{\,n}^{\,n+1}$, $T_{\,n}^{\,2}$ with $T_{\,n}^{\,n}$, and so on. Towers within each pair are constructed from identical-looking building blocks, but differ by a global rearrangement that leads to different enclosed cavity volumes. This pairing is made precise in the following corollary.
Corollary 2.
The conjugate tower pairs are given by $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$, where $i=1,2,\dots,n+1$. That is, a tower at fragmentation stage $i$ is paired with the tower at fragmentation stage ni+2n-i+2.
Proof.
If a tower at fragmentation stage $i$ is paired with a tower at fragmentation stage $p$ (without loss of generality, $i\leqslant p$ may be assumed), then the length $l_{i}$ of the building blocks at stage $i$ is equal to the height $h_{p}$ of the tower at the paired stage $p$. Using Eqs. (6) and (7), this condition can be written as
$2^{\,n-i+1}a=2^{\,p-1}a\,.$
From this relation, the paired fragmentation stage is obtained as p=ni+2p=n-i+2. Substituting back, the conjugate tower pairs are $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$, which completes the proof. ∎
Corollary 3.
The number of conjugate tower pairs generated from an initial object indexed by $n$ depends on the parity of $n$. For odd $n$, the $n+1$ towers form $(n+1)/2$ conjugate pairs. For even $n$, the towers form $n/2$ conjugate pairs, and the remaining tower $T_{\,n}^{\,n/2+1}$ is self-conjugate.
Proof.
The result follows directly from Corollary 2 by counting the pairs. ∎
Remark.
The emergence of conjugate and self-conjugate towers is a purely geometric consequence of the fragmentation–reassembly procedure. The term “geometric phase” is used here in an analogical sense, to denote distinct reassembled configurations of identical building blocks that enclose different volumes. No energetic or thermodynamic interpretation is implied.
III.4 General volume formula
Motivated by the explicit examples presented in the preceding subsections, a general expression for the maximal volume attained at each fragmentation stage is derived.
The volume of a tower is given by the sum of the material volume $2^{n+2}a^{3}$ of the initial parent object, which remains unchanged throughout the fragmentation process, and the volume of the enclosed space (cavity). The cavity volume is given by the product of the cross-sectional area of the cavity and the height of the tower, with the latter given by Eq. (7). The cross-sectional area of the cavity is equal to the square of the building-block length $l_{i}$, given by Eq. (6). The enclosed cavity volume is therefore expressed as
$l_{i}^{2}h_{i}=2^{\,2n-(i-1)}a^{3}.$
The total tower volume then follows immediately.
Proposition 2.
The volume $V_{\,n}^{\,i}$ of the tower $T_{\,n}^{\,i}$ reassembled at fragmentation stage $i$ from an initial parent object of dimensions $2^{\,n+2}a\times a\times a$ is given by
$V_{\,n}^{\,i}=2^{\,n+2}a^{3}+2^{\,2n-(i-1)}a^{3},\qquad i=1,2,\dots,n+1.$ (9)
The corresponding relative volume is obtained as
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}=\frac{2^{\,n+2}a^{3}+2^{\,2n-(i-1)}a^{3}}{2^{\,n+2}a^{3}}=1+2^{\,n-i-1}.$ (10)
The variation of the relative volume as a function of the fragmentation stage $i$ for several values of $n$ is illustrated in Fig. 4. As established in Corollary 1, the final fragmentation occurs at stage $i=n+1$. The relative volume at the final stage is therefore independent of $n$ and is given by
$R_{\,n}^{\,n+1}=1+2^{-2}=\frac{5}{4}.$ (11)
The value $5/4$ thus represents a geometric upper bound for the relative volume obtained after complete fragmentation within the present model.
The plots further indicate that the first fragmentation stage produces the largest increase in volume, after which the relative volume decreases monotonically with increasing stage index. This behaviour is governed by the evolution of the central cavity: the initial object encloses no void space, whereas the first fragmentation generates the largest possible cavity. At each subsequent stage, the cross-sectional area of the cavity is reduced by a factor of four, while the height of the tower increases only by a factor of two. Consequently, the enclosed cavity volume — and hence the cavity contribution to the relative volume — is reduced by a factor of two at each stage. The relative volume itself decreases monotonically but does not halve, owing to the additive contribution of the material volume.
[…]
Figure 4: Relative volume $R_{\,n}^{\,i}$ as a function of fragmentation stage $i$ for several values of $n$. Each curve corresponds to a fixed initial length $l=2^{n+2}a$. The first fragmentation stage yields the maximal relative volume, followed by a monotonic decrease towards the limiting value $R=5/4$ at $i=n+1$.
III.5 Remarks on paired configurations and “phase” interpretation
The pairing structure is closely related to the symmetry properties of the tower configurations.
The relative-volume difference for a conjugate pair (i,n+2i)(i,n+2-i) is given by
$\Delta R_{\,n}^{\,i}$ $=R_{\,n}^{\,i}-R_{\,n}^{\,n+2-i}=\bigl(1+2^{\,n-i-1}\bigr)-\bigl(1+2^{\,n-(n+2-i)-1}\bigr)$
$=2^{\,n-i-1}-2^{\,i-3}.$ (12)
This difference is maximal for $i=1$ (corresponding to the pairing of the first and last towers) and decreases as the index $i$ approaches the central value, $i=n/2+1$.
The relative-volume difference vanishes when
$i=\frac{n}{2}+1\,.$ (13)
Since $i$ must be a positive integer, this condition requires careful interpretation. If $n$ is even, Eq. (13) yields an integer value, and the pair (i,n+2i)(i,n+2-i) reduces to $(n/2+1,n/2+1)$. In this case, two conjugate towers coincide and form a single self-conjugate tower. If $n$ is odd, such a reduction does not occur, and a self-conjugate tower never exists.
Conjugate towers are characterized by a rearrangement of identical-looking building blocks. When these towers coincide in the even-$n$ case, the resulting self-conjugate tower inherits properties of both configurations. It can therefore be inferred that the self-conjugate tower may be constructed from the same building blocks in two distinct orientations, implying the presence of an internal symmetry, in agreement with the behaviour already observed in Sec. III.3.
These paired configurations provide a simple geometric analogue of distinct phases. While they are composed of identical-looking building blocks, they differ by a global rearrangement that modifies the enclosed cavity volume. Rearrangement-induced transitions between conjugate towers thus correspond to transitions between geometrically distinct configurations with different effective densities.
The geometric transitions observed in the present model admit a natural interpretation in the language of phase transitions. The relative volume $R_{\,n}^{\,i}$ plays the role of a macroscopic order parameter distinguishing tower configurations. Transitions between conjugate towers involve discontinuous changes in $R_{\,n}^{\,i}$ and may therefore be regarded as first-order-like.
In contrast, when $n$ is even, the middle tower is self-conjugate: rearrangement of the building blocks alters their internal orientation without changing the overall geometry or volume. This transition is associated with a discrete rotational symmetry and corresponds to a change in internal symmetry rather than in macroscopic structure. In this sense, the self-conjugate tower represents a geometric analogue of a second-order phase transition.
In summary, the fragmented tower configurations are organized into conjugate pairs distinguished by a discrete symmetry of relative volume, with a self-conjugate configuration arising only for even $n$. This structure provides a geometric analogue of phase behavior, in which conjugate towers represent distinct macroscopic phases, while the self-conjugate tower corresponds to a symmetry-driven transition without a change in density.
IV Connection to granular systems
To connect the present model with real granular systems, Eq. (10) is employed to evaluate the relative volume accessible to grains in experimental systems, while Eq. (12) is employed to identify the geometric phase transitions predicted by the model. Both expressions are formulated in terms of the discrete parameters $n$ and $i$. To enable experimental comparison, these parameters must be re-expressed in terms of measurable grain characteristics.
Throughout this section, it is assumed that grains do not interact with one another, that frictional effects are negligible, and that individual grains may be regarded as effectively weightless, such that no load is transmitted through the assembly. These assumptions are adopted in order to isolate the purely geometric contributions to volume evolution that are central to the present model.
IV.1 Remarks on experimental observation of the relative volume
A granular sample composed of identical grains with square cross-section of side length $a$ is considered. It is further assumed that these grains may be regarded as having been produced by sequential slicing of a hypothetical elongated parent object, in accordance with the fragmentation rule introduced earlier.
By comparing the total mass of the granular sample with the mass of an individual grain, the total number of grains in the sample, denoted by $N_{g}$, is determined. Let $l_{g}$ denote the length of an individual grain. The length $l$ of the hypothetical parent object is then given by
$l=N_{g}\,l_{g}\,.$ (14)
Within the model, the same length is expressed by Eq. (1) in terms of the discrete parameter $n$. By equating the two expressions and solving for $n$, one obtains
$n=\log_{2}\left(\frac{N_{g}\,l_{g}}{a}\right)-2\,.$ (15)
Here, the parameter $a$ is assigned the same meaning in both the model description and the experimental system.
The next step is to identify the fragmentation stage of the hypothetical parent object at which the size of a model grain coincides with that of the grains used in the experiment. At fragmentation stage $i$, the grain length $l_{i}$ (given by Eq. (6)) must match the experimental grain length $l_{g}$, so that
$2^{\,n-i+1}a=l_{g}\,.$
Using Eq. (15), this condition yields
$i=\log_{2}N_{g}-1\,.$ (16)
As a consistency check, for a single grain ($N_{g}=1$), Eq. (16) yields $i=0$, corresponding to the absence of fragmentation, as expected.
Substituting Eqs. (15) and (16) into Eq. (10) yields
$R=1+\frac{l_{g}}{4a}\,.$ (17)
This result demonstrates that the relative volume is determined exclusively by the geometric parameters of the grains and is independent of the details of the fragmentation history. All dependence on the hypothetical parent object is eliminated, rendering the indices $n$ and $i$ redundant; they are therefore omitted from Eq. (17).
Although the relative volume $R$ is expressed in Eq. (17) as a single function of the geometric ratio $l_{g}/a$, this quantity should be understood as a characteristic upper bound rather than as a uniquely reproducible experimental outcome. Under nominally identical preparation conditions, repeated realizations of the system are expected to yield a distribution of measured relative volumes clustered below this bound.
For a granular assembly composed of grains with square cross section, the relative volume may therefore be determined experimentally by measuring only the grain length $l_{g}$ and the side length $a$ of the cross section. Upon successive fragmentation, these measurements may be repeated until the terminal stage is reached, at which $l_{g}=a$. In this limit, the relative volume, given by Eq. (17), approaches the value $5/4$, in agreement with Eq. (11).
Exact quantitative agreement with the theoretical curve is not expected under experimental conditions. Nevertheless, the measured values are expected to follow the qualitative trend predicted by Eq. (17) and to remain bounded above by the maximal-volume curve provided by the model. Assuming square cross sections, these considerations may be summarised as follows.
Theorem (Liza’s theorem).
If a grain has length $l_{g}$ and a square cross section with side length $a$, the upper bound of the relative volume is determined solely by the ratio $l_{g}/a$ according to Eq. (17).
Remark (Liza’s limit).
The upper bound of the relative volume decreases monotonically as the grain length $l_{g}$ decreases. Its absolute minimum is attained when $l_{g}=a$, at which point the upper bound equals $5/4$, a value referred to as Liza’s limit or Liza’s number. The upper bound is independent of the number of grains and always exceeds unity, corresponding to the volume obtained by packing all grains adjacently without voids.
Equation (17) may be tested against existing experimental data. In Ref. [10], experiments were conducted on monodisperse nylon 6/6 cylindrical rods with diameter $1.8\,\mathrm{mm}$ and length $7.0\,\mathrm{mm}$. Approximately 7200 rods were poured into a vertical glass tube of height $1\,\mathrm{m}$ and inner diameter $1.90\,\mathrm{cm}$. Although the present model is formulated for grains with square cross section, a reasonable estimate is expected to be obtained for cylindrical rods of comparable transverse dimensions. The rods are therefore approximated as grains with square cross section of side length $a=1.8\,\mathrm{mm}$ and grain length $l_{g}=7.0\,\mathrm{mm}$. Substitution into Eq. (17) then yields
$R=1+\frac{l_{g}}{4a}\simeq 1.97\,.$
Experimentally, the quantity reported is the packing fraction $\rho$ (also commonly denoted $\phi$), defined as the fraction of the available volume occupied by the particles. When the longest grain dimension is small compared to the characteristic dimensions of the container — in the present case, the rod length relative to the tube diameter — the packing fraction may be approximated as $\rho\simeq R^{-1}$. This estimate gives $\rho\simeq 0.51$, which lies within the experimentally observed range $\rho=0.49$$0.55$ reported in Ref. [10].
IV.2 Comparison with experimental scaling laws
Experimental measurements have shown that the packing fraction $\phi$ of randomly oriented elongated particles decreases with increasing aspect ratio according to an asymptotic scaling law [8, 13]
$2\phi\times\alpha=\langle\gamma\rangle,\qquad\alpha\gg 1,$ (18)
where $\alpha=L/D$ is the aspect ratio, $L$ is the rod length, $D$ is the rod diameter, and $\langle\gamma\rangle$ is the average number of contacts per particle. Experiments indicate $\langle\gamma\rangle\approx 10.8$ for randomly packed rods [13].
The present geometric model can be compared with this scaling in the limit of large aspect ratio. For a grain with length $l_{g}$ and square cross-section of side $a$, the aspect ratio is $\alpha=l_{g}/a$. The packing fraction is related to the relative volume by $\phi=1/R$. Substituting Eq. (17) yields
$\phi=\frac{1}{R}=\frac{1}{1+l_{g}/(4a)}=\frac{4}{4+\alpha}.$ (19)
For large aspect ratios ($\alpha\gg 1$), this reduces to
$\phi\approx\frac{4}{\alpha},\qquad\text{or equivalently}\qquad 2\phi\times\alpha\approx 8.$ (20)
This result demonstrates that the present purely geometric model captures the correct $\phi\propto 1/\alpha$ scaling observed experimentally. The numerical coefficient differs from the experimental value ($8$ versus $\approx 10.8$), which is expected given that the model provides an upper bound on accessible volume and assumes maximal-volume configurations rather than typical random packings. The agreement in scaling behavior confirms that volume evolution under fragmentation, as described by the present model, is consistent with the experimental observation that packing densities are size-invariant and have a purely geometric origin [13].
Figure 5 compares the model predictions with experimental data from Freeman et al. [9] for aspect ratios ranging from $\alpha=4$ to $\alpha=32$. The experimental packing fractions lie consistently above the theoretical curve, as expected for a model that provides an upper bound on relative volume (or equivalently, a lower bound on packing fraction). The agreement in scaling behavior ($\phi\propto 1/\alpha$) validates the geometric approach.
$\alpha$$\phi$05101520253000.10.20.30.40.50.6Model: $\phi=4/(4+\alpha)$Experimental $\phi_{\infty}$Asymptotic: Eq. (18)
Figure 5: Experimental packing fraction $\phi_{\infty}$ as a function of aspect ratio $\alpha$ (circles with downward error bars) from Freeman et al. [9] compared with the present model (Eq. (19), solid blue line) and the asymptotic scaling law (Eq. (18), dashed blue line). The model provides a lower bound on packing fraction, while the asymptotic scaling law provides an empirical upper bound based on contact statistics.
The asymptotic scaling law (Eq. (18)), shown as a dashed line in Fig. 5, provides an empirical upper bound on packing fraction based on contact statistics in random packings. The experimental data lie between the two theoretical curves: above the geometric lower bound from the present model and below (or near) the empirical upper bound from the asymptotic scaling law. This positioning is physically reasonable, as real random packings achieve packing fractions intermediate between the maximal-volume geometric configurations considered here and the contact-limited configurations reflected in the asymptotic scaling. The convergence of both theoretical curves at large $\alpha$, with experimental data consistently confined between them, confirms that geometric packing constraints dominate throughout the range of aspect ratios studied, with the asymptotic regime exhibiting universal scaling behavior.
IV.3 Remarks on experimental observation of geometric phase transitions
An important question remains: can the geometric, rearrangement-induced phase transitions described above be observed experimentally? In addressing this question, two physical constraints must be taken into account. First, once a grain has been sliced, cut, or crushed, the resulting fragments cannot spontaneously reattach to reconstitute the original solid grain. Second, the levels of a tower are not mechanically bonded to one another; the upper levels merely rest on the lower ones under gravity.
A tower at fragmentation stage $i$ is constructed from grains of length $l_{i}=2^{\,n-i+1}a$ and has a height $h_{i}=2^{\,i-1}a$ (see Eqs. (6) and (7)). The conjugate tower at stage ni+2n-i+2 is instead constructed from grains of length $l_{n-i+2}=2^{\,i-1}a$ and has a height $h_{n-i+2}=2^{\,n-i+1}a$.
At first glance, the two towers appear to be assembled from identical grains: in the former case the grains are arranged horizontally, whereas in the latter they are aligned vertically. Crucially, however, this apparent equivalence is misleading. The vertically oriented “grain” is not a single object, but rather a stack of smaller fragments, whereas the horizontally oriented grain is an intact solid piece.
For illustration, the case $n=2$, shown in Fig. 3, may be considered. At fragmentation stage $i=1$, the grain length is $l_{1}=2^{\,2-1+1}a=4a$, and the tower consists of a single level of height $h_{1}=2^{\,1-1}a=a$. At the conjugate stage ni+2=3n-i+2=3, the grain length is $l_{3}=2^{\,i-1}a=a$, and the tower height is $h_{3}=2^{\,2-1+1}a=4a$. Four unit blocks are stacked vertically, producing an object that appears identical, at a coarse-grained geometric level, to a single grain of length $4a$ at stage $i=1$.
As discussed earlier, a tower at fragmentation stage $i$ is paired with a tower at stage ni+2n-i+2. This pairing reflects the fact that the two configurations are composed of building blocks that are geometrically indistinguishable at the level of their external shape, despite differing in their internal structure. Without loss of generality, the case i<ni+2i<n-i+2 is considered. In this situation, the grains at stage ni+2n-i+2 are smaller than those at stage $i$. Although their vertical alignment reproduces the geometry of the larger grains, recombination into the intact grains of stage $i$ is not possible. By contrast, the grains at stage $i$ can be rearranged to form the tower corresponding to stage ni+2n-i+2.
It is therefore concluded that the geometric phase transition is intrinsically directional and may occur only in the direction $i\to n-i+2$. However, if the intact grains at stage $i$ happen to be oriented vertically — matching the orientation of the stacked fragments at stage ni+2n-i+2 — then rotation to horizontal alignment enables the reverse transition $n-i+2\to i$ without requiring recombination of fragments.
With this constraint in mind, the following experimental situation may be considered. Suppose that $N_{g}$ grains, each of length $l_{g}$, are given. The fragmentation stage $i$ of the hypothetical parent object corresponding to this configuration is determined by Eq. (16). The conjugate stage $i_{c}\equiv n-i+2$, expressed directly in terms of experimentally measurable parameters, is then given by
$i_{c}=\log_{2}\!\left(\frac{l_{g}}{a}\right)+1.$ (21)
If
$i<i_{c},$ (22)
the occurrence of a geometric phase transition is permitted; otherwise, such a transition is prohibited.
By substituting the expressions for $n$ and $i$ from Eqs. (15) and (16) into Eq. (12) for the difference between the relative volumes of conjugate configurations, the following expression is obtained:
$\Delta R=\frac{l_{g}}{4a}-\frac{1}{16}N_{g}.$ (23)
Here, the indices on $\Delta R$ are redundant and have therefore been omitted.
As already assumed above, the comparison is made under the convention i<ni+2i<n-i+2, adopted without loss of generality. Since the relative volume decreases monotonically with increasing fragmentation stage, the difference $\Delta R$ is therefore non-negative. Equation (23) then yields the criterion
$\frac{N_{g}a}{4\,l_{g}}\leqslant 1.$ (24)
Equality corresponds to the self-conjugate configuration, in which the transition involves a change in internal geometric symmetry without any change in the enclosed volume.
The phase-transition criterion (24) provides insight into the interpretation of Eq. (17). When this criterion is not satisfied, the system admits a unique geometric configuration, and repeated experimental measurements of the relative volume $R$ are expected to form a single cluster centred around the characteristic value prescribed by Eq. (17). If the criterion is satisfied, however, two conjugate geometric configurations become accessible. In this regime, repeated preparations of the system may populate either configuration, resulting in two distinct clusters in the measured values of $R$, each corresponding to one of the conjugate geometric phases.
If condition (24) is not satisfied, the number of grains $N_{g}$ may, in principle, be reduced (provided $N_{g}\gg 1$) until the criterion is met. Combining these requirements yields
$1\ll N_{g}\leqslant\frac{4l_{g}}{a},$
indicating that the occurrence of a geometric phase transition requires grains with a sufficiently large aspect ratio.
The implications of this criterion may be illustrated by concrete examples. For the nylon rods discussed in Sec. IV.1 (with $l_{g}=7.0\,\mathrm{mm}$ and $a=1.8\,\mathrm{mm}$), the phase-transition condition implies $N_{g}<16$. As a further example with larger aspect ratio, matchsticks with tips removed, characterised by a cross-sectional side length $a=2\,\mathrm{mm}$ and length $l_{g}=5\,\mathrm{cm}$, yield the condition $N_{g}<100$.
In both cases, the number of grains required for the observation of the transition is relatively small. It is therefore concluded that geometric phase transitions of the type described here are restricted to mesoscopic systems, which may explain why they are not commonly observed in everyday granular materials.
Experimental support for the local nature of these transitions may be found in X-ray tomography studies. Freeman et al. [9] observed local alignment of small groups of rods (forming what may be termed domains) in containers filled with grains of aspect ratio $\alpha=8$. For these rods in containers of diameter $\approx 4.9\,l_{g}$, the criterion (24) requires $N_{g}<32$ for geometric phase transitions to occur. The observed local domains, while themselves randomly oriented within the container, are consistent with geometric rearrangements occurring in small grain clusters rather than uniformly throughout the sample. Similarly, Zakine et al. [5] found that pistachio shells — whose shapes deviate significantly from the idealised grains considered here — compact and interlock under mechanical agitation, suggesting that such transitions are facilitated by local rearrangements within small clusters.
This naturally motivates consideration of domains whose grain number and aspect ratio satisfy the phase-transition criterion derived above. Local rearrangements of this type may then cumulatively contribute to the macroscopic volume reduction observed during agitation. To test this hypothesis experimentally, the bulk packing density may be modelled as a superposition of contributions from different geometric configurations (or phases).
Although the present model is purely geometric and entirely static in nature, any experimental realisation necessarily involves real grains subject to friction, gravity, and contact constraints. Gentle mechanical agitation is therefore introduced not to endow the system with thermodynamic character, but simply to facilitate internal rearrangements and allow the system to explore its admissible geometric configurations. In this sense, agitation serves as a practical means of reducing kinetic trapping and history dependence, rather than as a source of thermalisation or energetic driving.
Experimental signatures.
The geometric phase transition mechanism admits two distinct experimental regimes depending on the total grain number.
Small-system regime ($N_{g}\lesssim 4l_{g}/a$): When the criterion (24) is satisfied for the entire grain assembly, repeated preparations under controlled agitation should yield a bimodal distribution of bulk relative volumes $R$, with peaks separated by $\Delta R$ (Eq. (23)) corresponding to global conjugate configurations.
Large-system regime ($N_{g}\gg 4l_{g}/a$): When the total grain number exceeds the criterion, global phase transitions are prohibited, but local transitions within domains containing $N_{g}\lesssim 4l_{g}/a$ grains remain possible. In this regime, the bulk relative volume forms a single distribution, but X-ray tomography should reveal aligned grain clusters (domains) of characteristic size $N_{\text{domain}}\sim 4l_{g}/a$. Since this constraint scales linearly with the aspect ratio $\alpha=l_{g}/a$, experiments conducted on grains of systematically varying aspect ratios should reveal that characteristic domain sizes increase proportionally with $\alpha$. Observation of such scaling would provide direct experimental evidence for the geometric phase transition mechanism, while its absence would indicate alternative origins for the local alignment patterns.
V Summary
A geometric model of granular matter composed of grains with square cross section has been developed by representing the system as originating from a hypothetical elongated parent object. The characteristic size of this object is fixed by the number and dimensions of the constituent grains.
Fragmentation is implemented through a well-defined slicing algorithm, applied iteratively and described in detail in the main text. At each stage of fragmentation, the resulting fragments are reassembled into configurations of maximal admissible volume, referred to as towers. The volumes of these towers were analysed as functions of the fragmentation stage. In realistic granular systems composed of elongated grains, the constituents at a given stage of fragmentation are expected to arrange in a largely unconstrained manner, resulting in occupied volumes that lie between those of compact packing and the largest geometrically admissible configurations. The construction of maximal-volume towers therefore provides a purely geometric upper bound on occupied volume, or equivalently, a lower bound on packing fraction, and allows the characteristic scale and qualitative trends of volume changes induced by fragmentation to be estimated.
It has been shown that, at the initial stage of fragmentation, the occupied volume increases relative to that of the parent object. With further fragmentation, the volume then decreases monotonically, approaching at the final stage the limiting value $5/4$ of the original volume. It follows that, throughout the fragmentation process, a material composed of elongated grains typically occupies a volume larger than that of the original object, unless the grains are packed perfectly without voids.
The inverse of the relative volume, corresponding to the packing fraction, was shown to obey the asymptotic scaling law $\phi\propto 1/\alpha$ for large aspect ratios, in agreement with experimental observations. Comparison with experimental data from Freeman et al. [9] demonstrates that the model provides a geometric lower bound on packing fraction, with experimental values lying consistently above the theoretical predictions, as expected for realistic random packings.
Within the fragmentation sequence, pairs of configurations have been identified and termed conjugate towers. These towers are composed of geometrically indistinguishable building blocks but enclose different volumes. When the number of fragmentation stages is even, a self-conjugate tower appears at the centre of the sequence. The magnitude of the volume difference between conjugate towers depends on their separation along the fragmentation sequence: towers corresponding to nearby stages exhibit smaller differences, whereas the largest difference occurs between the first and final stages of fragmentation.
Transitions between conjugate towers may be interpreted as geometric phase transitions. A criterion for the observability of such transitions has been derived, showing that they are restricted to small systems satisfying $N_{g}\lesssim 4l_{g}/a$. In large bulk systems where this criterion is not globally satisfied, geometric phase transitions may nevertheless occur locally within small domains of aligned grains. X-ray tomography studies of rod packings reveal such local alignment patterns, with characteristic domain sizes of order $N_{\text{domain}}\sim 4l_{g}/a$ grains, consistent with the theoretical prediction. A directly testable experimental signature is proposed: domain size should scale linearly with the aspect ratio $\alpha$, providing a quantitative test of the geometric phase transition mechanism.
In summary, an analytically tractable geometric model of fragmentation and reassembly has been presented, revealing features of volume evolution and phase-like behaviour in granular systems. Despite its simplicity, the model captures essential geometric aspects of granular matter, provides quantitative predictions in agreement with experimental scaling laws, and offers a framework for investigating packing geometry, porosity, metastable configurations, and local ordering phenomena in elongated grain assemblies.
Acknowledgements
The author thanks T. Börzsönyi for valuable comments on the role of Van der Waals forces in fine powder packing and for helpful discussions during the preparation of this manuscript.
The author is grateful to A. Zaccone for his interest in this work and for suggesting the review article Ref. [4].
The author also thanks E. R. Weeks for kindly providing numerical data from the experiments reported in Ref. [9].
This work is dedicated to the memory of the author’s grandmother, Liza Tsertsvadze. Many years ago, she remarked that a given mass of corn kernels occupies a smaller volume than the flour obtained from them. This simple observation provided the original motivation for the present study and inspired the investigation of the geometric origins underlying such volume changes.
The author acknowledges the use of AI language models (ChatGPT and Claude) to assist with manuscript preparation, including text editing, LaTeX formatting, and literature organization. All scientific content, theoretical development, and conclusions are solely the work of the author.